2NH4Cl(s)+Ba(OH)2⋅8H2O(s)→2NH3(aq)+BaCl2(aq)+10H2O(l)
The ΔH for this reaction is 54.8 kJ . How much energy would be absorbed if 26.1 g of NH4Cl reacts?
1
Expert's answer
2018-03-20T08:37:16-0400
2NH4Cl(s) + Ba(OH)2·8H2O(s) → 2NH3(aq) + BaCl2(aq) + 10H2O(l) , ∆H = 54.8 kJ

M(NH4Cl) = 14+4*1+35.5 = 53.5 g/mol

26.1g*54.8kJ / 2*53.5g = 13.37 kJ